OpenStudy (anonymous):

In 1980, the number of African elephants was 1,200,000 but that number was decreasing 5.25% a year. How many African elephants were there in 1988? Round your answer to the nearest whole number.

OpenStudy (ghostgate):

5.25% of 1,200,000 is 228,571.4 Now subtract: 1,200,000 - 228,571.4 = 971,428.6 = Elephants in 1981 5.25% of 971,428.6 is 185,034 Subtract: 971,428.6 - 185,034 = 786,394.6 = Elephants in 1982 5.25% of 786,394.6 is 149,789.5 Subtract: 786,394.6 - 149,789.5 = 636,605.1 = Elephants in 1983 5.25% of 636,605.1 is 121,258.1 Subtract: 636,605.1 - 121,258.1 = 515,347 = Elephants in 1984 5.25% of 515,347 is 98,161.3 Subtract: 515,347 - 98,161.3 = 417,185.7 = Elephants in 1985 5.25% of 417,185.7 is 79,463.9 Subtract: 417,185.7 - 79,463.9 = 337,721.8 = Elephants in 1986 5.25% of 337,721.8 is 64,328 Subtract: 337,721.8 - 64,328 = 273,393.8 = Elephants in 1987 5.25% of 273,393.8 is 52,075 Subtract: 273,393.8 - 52,075 = 221,318.8 = Elephants in 1988 If my math is correct than the answer your looking for is that: 221,319 African elephants remain in 1988.